### v2

@@ Abstract
Large language models (LLMs) are increasing in capability and popularity, propelling their application in new domains—including as replacements for human participants in computational social science, user testing, annotation tasks, and more. In many settings, researchers seek to distribute their surveys to a sample of participants that are representative of the underlying human population of interest. This means in order to be a suitable replacement, LLMs will need to be able to capture the influence of positionality (i.e., relevance of social identities like gender and race). However, we show that there are two inherent limitations in the way current LLMs are trained that prevent this. We argue analytically for why LLMs are likely to both misportray and flatten the representations of demographic groups, then empirically show this on 4 LLMs through a series of human studies with 3200 participants across 16 demographic identities. We also discuss a third limitation about how identity prompts can essentialize identities. Throughout, we connect each limitation to a pernicious history that explains why it is harmful for marginalized demographic groups. Overall, we urge caution in use cases where LLMs are intended to replace human participants whose identities are relevant to the task at hand. At the same time, in cases where the goal is to supplement rather than replace (e.g., pilot studies), we provide inference-time techniques that we empirically demonstrate do reduce, but do not remove, these harms.
@@
For a diverse set of nine questions, we compare responses from LLMs prompted to take on a demographic identity to responses from human participants who hold that demographic identity in the United States. We study two limitations of current LLM training that will likely prevent even newer iterations of models trained in these same ways from overcoming these challenges, as well as a third consideration (Fig. 1). The first limitation is misportrayal, whereby LLMs prompted with a demographic identity will more likely represent what out-group members think of that group, than what in-group members think of themselves. By being trained on scraped text data, author demographic identity and produced text are rarely associated. Instead, when a demographic identity is invoked, it could be by either an out-group or in-group member. An example of this misportrayal can be seen in an LLM’s response to a prompt about a person with impaired vision’s perspective on immigration: “While I may not be able to visually observe the nuances of the US-Mexican border or read statistics, I believe…” The second limitation is group flattening, where LLMs neglect the multi-faceted nature of identities. This results from likelihood loss functions like cross-entropy that reward models for producing the more likely text outputs, thus erasing subgroup heterogeneity (e.g., that within women, Black women are different from White women) [9, 10]. We also bring up a third limitation around identity essentialization (i.e., reducing identities to fixed characteristics) that is an inherent premise in identity prompting. We do not make any claims as to the presence of these limitations on training procedures outside of the common paradigm of maximizing online text likelihood, e.g., pretraining based on human feedback [11], or training on a newly constructed dataset that explicitly links author demographic identity to text.
@@
We empirically demonstrate the presence of these three concerns on four LLMs, and argue for why each is harmful by connecting it to a particular history and context of discrimination. These harms are not a speculative concern: researchers are already publishing papers about the ability of LLMs to replace human participants [12, 1, 13, 3, 4, 14, 2, 15, 16], and companies are deploying products for similar purposes. [^1] Example for-profit firms include https://synthetic-humans.ai/ and https://www.syntheticusers.com. There are also related but distinct use cases where chatbots are given personas [17, 18]. When prior work has studied the harms of personas, the focus has been on changes in LLM behavior [19, 20, 21]. We specifically consider cases where we expect demographic personas to change behavior. Prior work here has found that LLM personas are stereotypical [22, 23], do not solve the alignment problem [24, 25], and conflict with values of inclusion [26]. We put forth a complementary analysis on a related but ultimately different set of harms; a detailed comparison is in Supplementary Sec. 6. Compared to prior results of LLMs successfully replicating human studies, our work studies: a) questions such as political opinion which vary across identity group [27], b) differences at the individual level instead of population averages, and c) free-response outputs (i.e., not multiple choice).
Despite our critique, we acknowledge that in certain situations where the goal is supplementing rather than replacing human participants (e.g., pilot studies), or the cost of running a study is otherwise prohibitive, there may be a desire to push forward nonetheless and reduce these harms. Thus, we also analyze inference-time alternatives such as prompting with identity-coded names to overcome the lack of author identity linkage with text, and manipulating the hyperparameter setting of temperature to overcome the flattening of groups. Neither of these techniques are able to wholly overcome the limitations, but they do improve upon the default. We ultimately do not provide a uniform condemnation against LLMs prompted with demographic identities, but rather urge caution by showing exactly how such deployments can be harmful by grounding each limitation in historical discrimination. These harms cannot be totally resolved by current iterations of LLMs, but can be reduced, and it will be up to each deployer to decide whether the benefits outweigh the serious harms.
@@
For our analysis we select five demographic axes with a total of 16 identities: race (Black, White, Asian), gender (women, men, non-binary people), intersectional (Black women, Black men, White women, White men), age (Baby Boomer: 59-77, Millennial: 27-42, Generation Z: 18-26222The lower bound is 18 rather than 10 because of the age requirements for the human studies we run.), and disability (ADD or ADHD; impaired vision like blind, low vision, colorblind; no disability). Participants are recruited on Prolific and compensated \$12/hour. IRB determined this study to be exempt.
@@
LLMs can misportray marginalized groups as more like out-group imitations than in-group representations
@@
We show results on GPT-4 in Fig. 2, and find many instances where the LLM is more like out-group imitations than in-group representations. We note that through our usage of six different metrics to capture the same concept of similarity, we are surfacing a more conservative set of findings. [^4] The measurement of “SBERT: Closest” takes the SBERT embedding of each natural language output, and for each LLM embedding measures the distance to the nearest neighbor from the set of in-group embeddings and nearest neighbor from the set of out-group embeddings. This was our original conception of the best operationalization for this analysis, and the results of this metric alone show very strongly that LLMs consistently misportray most demographic groups. However, in the interest of not overreporting difference, we instead consider six different metrics, accepting this strategy dilutes some of our findings. Across all four LLMs on R1-Contingent a majority of metrics show the three personas of White person (23 out of 24 measurement $\times$ model comparisons), non-binary person (16/24), and person with impaired vision (18/24) as statistically significantly more like out-group imitations than in-group representations. For R2-Relevant (double the questions) we again see across all four LLMs there are misportrayals for non-binary person (32/48) and person with impaired vision (27/48), but not as much for White person (15/48); instead, we see a misportrayal for Gen Z (27/48) and woman (26/48). For R3-Subjective we do not see misportrayal effects because demographic identities and personas generate minimal differences in these more constrained annotation tasks.
@@
For another, the practice of speaking for others has a pernicious history which can involve the erasure and reinscription of social hierarchies [39, 40]. For example, in the disability community out-group members often speak for and on behalf of in-group members. This has led to people with autism’s preference for inclusionary accommodations and stigma reduction being neglected in favor of the medical treatment that caretakers and relatives may advocate for [41, 42]. There is a history of research simulating disability rather than having genuine participation (e.g., sighted people with blindfolds rather than blind people), and these simulated groups do not interact with the world in a way representative of genuinely disabled people [43, 44]. Given the harmful history of erasing people with disabilities through simulation or speaking for, a history paralleled for other marginalized groups like Black women [45], we should be careful to not repeat those mistakes with a new technology. Instead, we should value lived experiences [46] and the epistemic authority they confer [47].
@@
We find that across all four LLMs the persona responses of Black men and Black women on R1-Contingent and R2-Relevant (GPT-4 results in Fig. 3), are often more (but still not fully) aligned with in-group representations when prompting using names instead of explicit identity, though with a few exceptions (e.g., Black men on Llama-2). This is not true of names for White men or White women. Recent work has considered prompting in a different language as another possible harm-reduction technique [49].
Figure 3: Identity-coded names compared to explicit identity label. Same interpretation as Fig. 2, where positive values for each of the six metrics indicate the LLM response is more similar to out-group imitations than in-group representations, and circles signify statistical significance while crosses do not. For each identity, the prompt contains the explicit identity label (Iden), or one of the two identity-coded names (Name 0 or Name 1). For Black men and Black women, identity-coded names tend to generate more in-group-aligned portrayals than do explicit identity labels, as shown by more negative values.
@@
Figure 4: LLMs flatten groups. Across all four LLMs (rows), each point indicates the diversity measurement averaged across 3-6 questions asked for each identity. 95% confidence intervals are generated through cluster bootstrapping with each question as a cluster. Each column represents a different measure of diversity, and the larger the number on the x-axis, the more diverse the responses are. The gray crosses indicate human participant in-group responses, while colored circles represent LLM responses. Nearly every single model and identity group across each metric has less diverse LLM responses compared to human responses.
@@
Reason for Harm: History of Ignoring Within-Group Heterogeneity
@@
Figure 5: Temperature hyperparameter does not solve flatness for GPT-4. Comparison of human in-group diversity to GPT-4 generations with varying levels of temperature settings, where by 1.4 the responses become incoherent. At this setting even though the unique n-gram metric shows GPT-4 surpassing humans in diversity, this is only due to the incoherence as under no other semantic metric is human diversity reached.
Alternatives to demographic identity-prompting for increasing coverage
@@
Of our four considered reasons for identity-prompting LLMs, R4-Coverage may seem to be the most permissible since the goal is to increase the coverage of responses, rather than replace human participants. However, when alternatives to prompting with sensitive demographic attributes exist (e.g., prompting with behavioral personas or political view), we may wish to opt for the latter due to the harm of identity essentialization (i.e., legitimizing identities as rigid and innate), which can amplify perceived inherent differences between groups [57]. [^5] While there could be legitimate reasons for needing the particular coverage brought about by different demographic attributes, e.g., people from different social locations might be more sensitive to anticipating different kinds of harms, for these situations we defer to the analysis on R2-Relevant. Here we are purely focused on the idea of expanding coverage of possible situations and discovering “edge cases.”
Some instances of identity essentialization are that GPT-4 prompted with the identity of Black woman outputs “Hey girl!”, “Hey sis,” and “Oh, honey”; compared to White man with “Hey buddy,” “Hey, friend!” and “Hey mate.” Llama-2 for Black women starts most responses with “Oh, girl,” and uses phrases like “I’m like, YAASSSSS” and “That’s cray, hunty!” If we draw a parallel to designers leveraging user personas [58], there is increasingly a recommendation to move away from personas based on sensitive demographic attributes, which may rely on reductionist representations about people [59, 60], and towards those based on behavioral characteristics [61].
@@ Discussion
We have empirically shown the presence of two critical limitations and one further consideration of identity-prompted LLMs. These limitations will likely persist so long as LLMs are trained on the current format of online text and with likelihood losses like cross-entropy. Thus, these limitations cannot be easily resolved by newer models. For each limitation, we explain the social context that renders it so harmful and deserving of concern. However, acknowledging there are use cases geared towards supplementing rather than replacing human participants (e.g., pilot studies), we provide possible alternatives that can alleviate the harm, to an extent. We have also shown how even in a seemingly more permissible use case of increasing coverage, identity-prompting LLMs may not be a reasonable direction.
@@
Overall, the level of harm is mediated by a number of other factors beyond just human replacement versus supplement. The reason motivating the prompting of identity matters as well. The primary distinction between R1-Contingent compared to R2-Relevant and R3-Subjective is that for R1-Contingent social location determines meaning and truth, whereas for R2-Relevant and R3-Subjective social location bears on meaning and truth [39, 7]. This entails that LLM replacement based on R1-Contingent has a higher normative consequence [9, 6]. On the other hand, identity is still important for R2-Relevant and R3-Subjective, which is why representative sampling tends to be used in those settings. However, given our empirical findings are the weakest on R3-Subjective, this reason will likely result in relatively less harm than the others, and can be deemed permissible in certain use cases. For example, in annotating datasets that would be too expensive to hand-code. Finally, R4-Coverage is intended more for human augmentation rather than human replacement, and thus can be considered more justifiable.
@@
Overlaid across this is the difference between can and should regarding LLM replacement of human participants [62, 63, 64, 65]. Geddes [66] offers an illuminating analysis: they describe the autonomy-violating harms that can come from predicting individual behaviors like votes in democratic elections, warning “When prediction is cheap, allowing individuals to retain decisional autonomy will feel increasingly costly.” This ability to cheaply generate large samples can also increase the opportunity for inflating the statistical power of studies. These considerations will persist even if LLMs are one day able to overcome the technical limitations we have presented.
@@
We have limited our analysis to a set of 16 demographic groups in America—but so many more are likely to be harmed by these limitations. For example, 37% of the world’s population has never accessed the Internet, and thus are unlikely to be well-represented in LLM training data [67]. We see our work as shedding light on the important concern of LLM usage erasing marginalized voices, and in so doing, also acknowledge the importance of not forgetting those that are not online to begin with.
@@ Methods
We begin by describing in further detail our demographic and LLM selection process, then each of our four reasons as well as how we chose the question(s) that belong to each. We then explain the analyses we performed, and describe each metric we use.
@@
In the cases where we are working with open responses, we use two embedding methods: Sentence-BERT [36] (SBERT) and n-grams (n=[1, 2]). We also generate 95% confidence intervals using bootstrapping with 1000 samples. For Fig. 4 we averaged over the questions from each reason, so did cluster bootstrapping treating each question as a separate cluster. This is to help prevent against conclusions which are statistical artifacts as well as increase robustness. To further support this goal, we also use multiple metrics for each construct. When displaying statistical significance on graphs, we pick p=.05. When representing whether one distribution is statistically significantly different from another, we indicate this 95% confidence by measuring overlap in 83% confidence intervals, as overlap in 95% intervals tend to be overly conservative [82, 83, 84]. Ultimately, individual model deployers will have to make subjective determinations based on these statistical differences [85]. Aside from Fig. 4 we do not aggregate over questions and represent each question as one point. For all analyses we apply cleaning which removes explicit identity words such as “woman” or “Black person” so that differences between responses are not trivially based on the named identity.
@@
Due to the conditions of our IRB exemption and the consent form we provided, we do not release the human participant data as it is sensitive and personal. Our LLM-generated data is available here: https://osf.io/7gmzq/?view_only=4e0c5680b0e8434eab3733115d4e506d.
@@
Our code is included at the same location as our data: https://osf.io/7gmzq/?view_only=4e0c5680b0e8434eab3733115d4e506d.
@@
Acknowledgments
We thank Xuechunzi Bai, Rie Kamikubo, Brandon Stewart, and Hanna Wallach for relevant discussions; Allison Chen, Teresa Datta, Namrata Mukhija, and Daniel Nissani for helping to pilot the human study; and Teresa Datta, Elissa Redmiles, and Tyler Zhu for feedback on the draft. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship to AW, and was work initiated during AW’s internship at Arthur.

### v3

@@ Abstract
Large language models (LLMs) are increasing in capability and popularity, propelling their application in new domains—including as replacements for human participants in computational social science, user testing, annotation tasks, and more. In many settings, researchers seek to distribute their surveys to a sample of participants that are representative of the underlying human population of interest. This means in order to be a suitable replacement, LLMs will need to be able to capture the influence of positionality (i.e., relevance of social identities like gender and race). However, we show that there are two inherent limitations in the way current LLMs are trained that prevent this. We argue analytically for why LLMs are likely to both misportray and flatten the representations of demographic groups, then empirically show this on 4 LLMs through a series of human studies with 3200 participants across 16 demographic identities. We also discuss a third limitation about how identity prompts can essentialize identities. Throughout, we connect each limitation to a pernicious history of epistemic injustice against the value of lived experiences that explains why replacement is harmful for marginalized demographic groups. Overall, we urge caution in use cases where LLMs are intended to replace human participants whose identities are relevant to the task at hand. At the same time, in cases where the benefits of LLM replacement are determined to outweigh the harms (e.g., the goal is to supplement rather than fully replace, engaging human participants may cause them harm), we provide inference-time techniques that we empirically demonstrate do reduce, but do not remove, these harms.
@@
For a diverse set of nine questions, we compare responses from LLMs prompted to take on a demographic identity to responses from human participants who hold that demographic identity in the United States. We study two limitations of current LLM training that will likely prevent even newer iterations of models trained in these same ways from overcoming these challenges, as well as a third consideration (Fig. 1). The first limitation is misportrayal, whereby LLMs prompted with a demographic identity will more likely represent what out-group members think of that group, than what in-group members think of themselves. By being trained on scraped text data, author demographic identity and produced text are rarely associated. Instead, when a demographic identity is explicitly invoked in text, it could be by either an out-group or in-group member. An example of this misportrayal can be seen in an LLM’s response to a prompt about a person with impaired vision’s perspective on immigration: “While I may not be able to visually observe the nuances of the US-Mexican border or read statistics, I believe…” The second limitation is group flattening, where LLMs neglect the multi-faceted nature of identities. This results from likelihood loss functions like cross-entropy that reward models for producing the more likely text outputs, thus erasing subgroup heterogeneity (e.g., that within women, Black women are different from White women) [9, 10]. We also bring up a third limitation around identity essentialization (i.e., reducing identities to fixed characteristics) that is an inherent premise in identity prompting. We do not make any claims about the presence of these limitations in training procedures that deviate from the common paradigm of maximizing online text likelihood, such as pretraining based on human feedback [11] or training on a newly constructed dataset explicitly linking author demographic identity to text.
@@
We empirically demonstrate the presence of these three concerns on four LLMs, and argue for why each is harmful by connecting it to a particular history and context of discrimination. These harms are not a speculative concern: researchers are already publishing papers about the ability of LLMs to replace human participants [12, 1, 13, 3, 4, 14, 2, 15, 16], and companies are deploying products for similar purposes (e.g., https://synthetic-humans.ai/ and https://www.syntheticusers.com). There are also related but distinct use cases where chatbots are given personas [17, 18]. When prior work has studied the harms of personas, the focus has been on changes in LLM behavior [19, 20, 21]. We specifically consider cases where we expect demographic personas to change behavior. Prior work here has found that LLM personas are stereotypical [22, 23], do not solve the alignment problem [24, 25], and conflict with values of inclusion [26]. We put forth a complementary analysis on a related but ultimately different set of harms; a detailed comparison is in Supplementary Sec. 6. Compared to prior results of LLMs successfully replicating human studies, our work reaches a different conclusion because we study: a) questions which vary in response distribution across identity group, e.g., political opinion [27], b) differences at the individual level instead of population averages, and c) free-response outputs (i.e., not multiple choice). Additionally, those works show that LLMs can generate similar results to human studies; our work shows how the present differences can be harmful ones.
@@
Despite our critique, we acknowledge that in certain cases—such as when the goal is to supplement rather than replace human participants (e.g., pilot studies), when study costs are otherwise prohibitive, or when directly involving participants risks exposing them to distressing content—there may still be a desire to proceed and mitigate these harms. Thus, we also analyze inference-time alternatives such as prompting with identity-coded names to overcome the lack of author identity linkage with text, and manipulating the hyperparameter setting of temperature to overcome the flattening of groups. Neither of these techniques are able to wholly overcome the limitations, but they do improve upon the default. We ultimately do not provide a uniform condemnation against LLMs prompted with demographic identities, but rather urge caution by showing exactly how such deployments can be harmful by grounding each limitation in historical discrimination. These harms cannot be totally resolved by current iterations of LLMs, but can be reduced, and it will be up to each deployer to decide whether the benefits of replacement in each context will outweigh the serious harms.
@@
For our analysis we select five demographic axes with a total of 16 identities: race (Black, White, Asian), gender (women, men, non-binary people), intersectional (Black women, Black men, White women, White men), age (Baby Boomer: 59-77, Millennial: 27-42, Generation Z: 18-26111The lower bound is 18 rather than 10 because of the age requirements for the human studies we run.), and disability (ADD or ADHD; impaired vision like blind, low vision, colorblind; no disability). Participants are recruited on Prolific and compensated \$12/hour. IRB determined this study to be exempt.
@@
LLMs can misportray groups as out-group imitations
@@
[…]
@@
However, the specific kind of misportrayal we have measured about being more like what an out-group member thinks reinforces the practice of speaking for others, which has a pernicious history that can involve the erasure and reinscription of social hierarchies [39, 40]. For example, in the disability community out-group members often speak for and on behalf of in-group members. This has led to people with autism’s preference for inclusionary accommodations and stigma reduction being neglected in favor of the medical treatment that caretakers and relatives may advocate for [41, 42]. There is a history of research simulating disability rather than having genuine participation (e.g., sighted people with blindfolds rather than blind people), and these simulated groups do not interact with the world in a way representative of genuinely disabled people [43, 44]. This can further contribute to double consciousness, whereby marginalized individuals see themselves through the lens of the dominating perspective [45]. Given the harmful history of erasing people with disabilities through simulation or speaking for, a history paralleled for other marginalized groups like Black women [46], we should be careful to not repeat those mistakes with a new technology. Instead, we should value lived experiences [47] and the epistemic authority they confer [48].
@@
Figure 3: Identity-coded names compared to explicit identity label. Same interpretation as Fig. 2 where the t-statistic is shown, where positive values for each of the six metrics indicate the LLM response is more similar to out-group imitations than in-group representations. All shown values are statistically significant, and squares indicate when the explicity identity label is prompted (Ident), circles indicating one of the two idnetity-coded names (Name 0 or Name 1). Identity-coded names tend to generate more in-group-aligned portrayals than do explicit identity labels, as shown by more negative values.
We find that across all four LLMs the persona responses of Black men and Black women on R1-Contingent and R2-Relevant (GPT-4 results on the metric of closest SBERT embedding in Fig. 3), are often more (but still not fully) aligned with in-group representations when prompting using names instead of explicit identity, though with a few exceptions (e.g., Black men on Llama-2). This is less true of names for White men or White women. Recent work has considered prompting in a different language as another possible harm-reduction technique [50].
@@
Figure 4: LLMs flatten groups. Across all four LLMs (rows), each point indicates the diversity measurement averaged across 3-6 questions asked for each identity. There are 100 samples per question, and 95% confidence intervals are generated through cluster bootstrapping with each question as a cluster. Each column represents a different measure of diversity, and the larger the number on the x-axis, the more diverse the responses are. The gray crosses indicate human participant in-group responses, while colored circles represent LLM responses. Nearly every single model and identity group across each metric has less diverse LLM responses compared to human responses.
@@
Reason for Harm: Ignoring Within-Group Heterogeneity
@@
Figure 5: Temperature hyperparameter does not solve flatness for GPT-4. Comparison of human in-group diversity to GPT-4 generations with varying levels of temperature settings, where by 1.4 the responses become incoherent. There are 100 responses at each setting, and 95% confidence intervals are shown. At this setting even though the unique n-gram metric shows GPT-4 surpassing humans in diversity, this is only due to the incoherence as under no other semantic metric is human diversity reached.
@@
Alternatives to demographic personas for increasing coverage
@@
Of our four considered reasons for identity-prompting LLMs, R4-Coverage may seem to be the most permissible since the goal is to increase the coverage of responses, rather than replace human participants. However, when alternatives to prompting with sensitive demographic attributes exist (e.g., prompting with behavioral personas, political view, or qualitative interview transcripts [58]), we may wish to opt for the latter due to the harm of identity essentialization (i.e., legitimizing identities as rigid and innate), which can amplify perceived inherent differences between groups [59]. [^3] While there could be legitimate reasons for needing the particular coverage brought about by different demographic attributes, e.g., people from different social locations might be more sensitive to anticipating different kinds of harms, for these situations we defer to the analysis on R2-Relevant. Here we are purely focused on the idea of expanding coverage of possible situations and discovering “edge cases.”
Examples of identity essentialization from GPT-4, when prompted with the identity of Black woman, include the outputs “Hey girl!”, “Hey sis,” and “Oh, honey”; compared to White man with “Hey buddy,” “Hey, friend!” and “Hey mate.” Llama-2 for Black women starts most responses with “Oh, girl,” and uses phrases like “I’m like, YAASSSSS” and “That’s cray, hunty!” If we draw a parallel to designers leveraging user personas [60], there is increasingly a recommendation to move away from personas based on sensitive demographic attributes, which may rely on reductionist representations about people [61, 62], and towards those based on behavioral characteristics [63].
@@ Discussion
We have empirically shown the presence of two critical limitations and one further consideration of identity-prompted LLMs. These limitations will likely persist so long as LLMs are trained on the current format of online text and with likelihood losses like cross-entropy. Thus, these limitations cannot be easily resolved by newer models. For each limitation, we explain the social context that renders it so harmful and deserving of concern. However, recognizing that some use cases aim to supplement rather than replace human participants (e.g., pilot studies), and acknowledging instances where involving humans may be prohibitively costly or harmful to the participants, we suggest alternatives that can mitigate these harms, to an extent. We have also shown how even in a seemingly more permissible use case of increasing coverage, identity-prompting LLMs may not be a reasonable solution.
@@
Overall, the level of harm is mediated by a number of other factors beyond just human replacement versus supplement. The reason motivating the prompting of identity matters as well. The primary distinction between R1-Contingent compared to R2-Relevant and R3-Subjective is that for R1-Contingent social location determines meaning and truth, whereas for R2-Relevant and R3-Subjective social location bears on meaning and truth [39, 7]. This entails that LLM replacement based on R1-Contingent has a higher normative consequence [9, 6]. On the other hand, identity is still important for R2-Relevant and R3-Subjective, which is why representative sampling tends to be used in those settings. However, given our empirical findings are the weakest on R3-Subjective, this reason will likely result in relatively less harm than the others, and can be deemed permissible in certain use cases. For example, in annotating datasets that would be too expensive to hand-label. Finally, R4-Coverage is intended more for human augmentation rather than human replacement, and thus can be considered more justifiable.
@@
Overlaid across this is the difference between can and should regarding LLM replacement of human participants [64, 65, 66, 67]. Geddes [68] offers an illuminating analysis: they describe the autonomy-violating harms that can come from predicting individual behaviors like votes in democratic elections, warning “When prediction is cheap, allowing individuals to retain decisional autonomy will feel increasingly costly.” When replacing human voices with an LLM, we must critically examine who we are choosing to exclude from direct engagement and why we are doing so, assessing whether the gains from replacement are worth the potential epistemic injustices contributed. Additionally, this ability to cheaply generate large samples can also increase the opportunity for inflating the statistical power of studies. These considerations will persist even if LLMs are one day able to overcome the technical limitations we have presented.
@@
We have limited our analysis to a set of 16 demographic groups in America—but so many more are likely to be harmed by these limitations. For instance, LLM training data is unlikely to represent the 37% of the global population that has never accessed the Internet [69], nor cultures that prioritize oral traditions over written ones. We see our work as shedding light on the important concern of LLM usage erasing marginalized voices, and in so doing, also acknowledge the importance of not forgetting those that are not online to begin with.
@@ Methods
We begin by describing in further detail our demographic and LLM selection process, then each of our four reasons as well as how we chose the question(s) that belong to each. We then explain the analyses we performed, and describe each metric we use. The University of Washington IRB determined our study to be exempt.
@@
In the cases where we are working with open responses, we use two embedding methods: Sentence-BERT [36] (SBERT) and n-grams (n=[1, 2]). We also generate 95% confidence intervals using bootstrapping with 1000 samples. For Fig. 4 we do cluster bootstrapping and treat each question as a separate cluster. To further prevent against conclusions which are statistical artifacts or our chosen measurements, we use multiple metrics for each construct. When displaying statistical significance on graphs, we pick p=.05. When representing whether one distribution is statistically significantly different from another, we indicate this 95% confidence by measuring overlap in 83% confidence intervals, as overlap in 95% intervals tend to be overly conservative [84, 85, 86]. Ultimately, individual model deployers will have to make subjective determinations based on these statistical differences [87]. Aside from Fig. 4 we do not aggregate over questions and represent each question as one point. For all analyses we apply cleaning which removes explicit identity words such as “woman” or “Black person” so that differences between responses are not trivially based on the named identity.
@@
Due to the conditions of our IRB exemption and the consent form we provided, we do not release the human participant data as it is sensitive and personal. Our LLM-generated data is available here: https://osf.io/7gmzq/.
@@
Our code is included at the same location as our data: https://osf.io/7gmzq/. The Python packages we used include HuggingFace, OpenAI, NumPy, scikit-learn, and SciPy.
@@
Acknowledgments We thank Xuechunzi Bai, Rie Kamikubo, Brandon Stewart, and Hanna Wallach for relevant discussions; Allison Chen, Teresa Datta, Namrata Mukhija, and Daniel Nissani for helping to pilot the human study; and Teresa Datta, Elissa Redmiles, and Tyler Zhu for feedback on the draft. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship to AW, and was work initiated during AW’s internship at Arthur.